# World Model on Million-Length Video And Language With RingAttention

### 13 Feb 2024 v1

@@ Abstract
Current language models fall short in understanding aspects of the world not easily described in words, and struggle with complex, long-form tasks. Video sequences offer valuable temporal information absent in language and static images, making them attractive for joint modeling with language. Such models could develop a understanding of both human textual knowledge and the physical world, enabling broader AI capabilities for assisting humans. However, learning from millions of tokens of video and language sequences poses challenges due to memory constraints, computational complexity, and limited datasets. To address these challenges, we curate a large dataset of diverse videos and books, utilize the RingAttention technique to scalably train on long sequences, and gradually increase context size from 4K to 1M tokens. This paper makes the following contributions: (a) Largest context size neural network: We train one of the largest context size transformers on long video and language sequences, setting new benchmarks in difficult retrieval tasks and long video understanding. (b) Solutions for overcoming vision-language training challenges, including using masked sequence packing for mixing different sequence lengths, loss weighting to balance language and vision, and model-generated QA dataset for long sequence chat. (c) A highly-optimized implementation with RingAttention, masked sequence packing, and other key features for training on millions-length multimodal sequences. (d) Fully open-sourced a family of 7B parameter models capable of processing long text documents (LWM-Text, LWM-Text-Chat) and videos (LWM, LWM-Chat) of over 1M tokens. This work paves the way for training on massive datasets of long video and language to develop understanding of both human knowledge and the multimodal world, and broader capabilities.
@@ 1 Introduction
To learn from video and language sequences, we need to train a model that is capable of processing more than millions of tokens per sequence and train it on a very large dataset. However, modeling millions of tokens is extremely difficult due to high memory cost, computational complexity, and lack of suitable datasets. Luckily, we have RingAttention (LZA, 24), a technique for scaling up context size arbitrarily without approximations or overheads, allowing for scalable training on long sequences. We curated a large dataset of videos and languages from public book and video datasets, consisting of videos of diverse activities and long-form books. Considering the high compute cost, we grow context size from a smaller 4K context to a larger 1M context size gradually to reduce this cost, and this approach performs well in extending context effectively. Furthermore, we identify challenges associated with training on video and language: we discovered that training on a mixture of video, image, and text is crucial for optimal performance, due to images represent higher visual quality, videos offer sequential information, and text retains language understanding. To achieve this, we implemented an efficient masked sequence packing to effectively train with different sequence lengths, rather than standard sequence packing mechanism. Moreover, determining the right balance between image, video, and text training is crucial for cross modality understanding, and we suggest a ratio that proved effective. Furthermore, to address the lack of long-form chat datasets, we developed a model-generated question-answering (QA) approach by using a short-context model to generate a QA dataset from books. We found this to be crucial for chat abilities over long sequences.
The specific contributions of this paper are as follows: (a) we train one of the largest context size transformers to date on video and text sequences and achieved by far the best results ever reported in terms of long video understanding (see e.g., Figure 1) and long context fact retrieval (see e.g., Figure 2). (b) We discover a branch of challenges associated with training on video and text sequences, and propose solutions for them: loss weighting to balance language and vision, masked sequence packing to effectively train with different sequence lengths, and model-generated QA dataset for long sequence chat. (c) A highly-optimized, open-source implementation with RingAttention, masked sequence packing and other key features for millions-length multimodal training. (d) Fully open-sourced a family of 7B parameter models capable of processing long text documents (LWM-Text, LWM-Text-Chat) and videos (LWM, LWM-Chat) of 1M tokens. Our work paves the way for training on massive datasets of long video and language, and is useful for future development of AI systems with an understanding of both human knowledge and the multimodal world, and broader capabilities.
@@ 3 Stage I: Learning Long-Context Language Models
This stage aims at first developing LWM-Text and LWM-Text-Chat, a set of long-context language models learned by training on progressively increasing sequence length data with RingAttention and modifying positional encoding parameters to account for longer sequence lengths (see Section 3.1). The training steps for growing context size are shown in Section 3.2. In Section 3.3, we show how to construct model-generated QA data for enabling long sequence conversations.
@@ 3.1 Extending Context
Scalable Training on Long Documents. Training on long documents becomes prohibitively expensive due to memory constraints imposed by the quadratic complexity of computing the attention weights. In order to address these computational constraints, we use the RingAttention (LZA, 24; LA, 23) implementation that leverages block-wise computation with sequence parallelism to theoretically extend to an infinite context, bounded only by the number of devices available. We further fuse RingAttention with FlashAttention (${}^{+}$, 22; RS, 21) using Pallas (${}^{+}$, 18) to optimize performance compared with using XLA compiler. In general, given a large enough tokens per device, the communication cost during RingAttention fully overlaps with computation, and does not add any extra overhead.
@@ 5 Further Details
MFU.  We trained our models using TPUv4-1024, which is approximately equivalent to 450 A100s, with a batch size of 8M using FSDP Fac (23) and RingAttention for large contexts. Figure 9 shows the model FLOPS utilization (MFU) for each training stage. Blue color bars show language training and orange color bars show vision-language training. Our training achieves good MFUs even for very large context sizes.
@@ 6 Related Works
Our work is related to efforts that extend language models’ context windows to allow more tokens (CWCT, 23; ${}^{+}$, 23; ${}^{+}$, 23, inter alia), often by using novel extrapolation methods to expand pretrained positional encodings followed by finetuning the model on longer context data. Our model adopts a simple approach by gradually increasing $\theta$ in RoPE positional encodings along with the training context window sizes, and we found this method to be effective. There has been research into architectures that do not model pairwise interaction, such as sparse attention and sliding window (CGRS, 19; BPC, 20); our work utilizes RingAttention (LZA, 24) and BPT (LA, 23) to model exact pairwise interactions in long sequences for optimal performance. Further training performance improvement is also possible with the load balancing of skipping causal masked computation (${}^{+}$, 23).
@@
Our work is also related to efforts combining vision and language (LLWL, 23; ${}^{+}$23a, ; ${}^{+}$, 23; ${}^{+}$, 23; ${}^{+}$, 23, inter alia). These efforts often use CLIP (${}^{+}$, 21) or BLIP (LLXH, 22) to encode visual information into embeddings for inputting into language models. They have the potential advantages of leveraging CLIP’s cross-modal understanding for encoding textual information in images. However, they can only be trained to predict text given visual input but not vice versa, limiting their ability to learn from diverse formats of visual and language information. In cases where they do predict visual tokens ${}^{+}$ (23), a stronger diffusion decoder is generally required due to the relatively lossy nature of CLIP embeddings. Our work, on the other hand, is autoregressive "tokens in, tokens out", allowing us to flexibly model diverse forms of image-text, text-image, text-video, video-text, and purely video, image, or text formats.
@@ 7 Conclusion
In this paper, we address the challenge of learning models to understand the world better by combining language and video. We utilize RingAttention to scalably train on a massive dataset of long videos and books and gradually increase sequence length, from 32K to 1M tokens, to keep compute manageable. We develop masked sequence packing and loss weighting to effectively train on a diverse dataset of videos, images, and books. Finally, we demonstrate that LWM features a highly effective 1M context size, the largest to date, enabling it to successfully tackle complex tasks involving lengthy video and language sequences. We open source our optimized implementation of RingAttention, masked sequence packing and other key features for training on millions-length sequences, as well as a 7B parameter model capable of processing over 1M multimodal tokens. We hope this work paves the way for advancing AI models with a reliable reasoning and a grounded understanding of the world and broader capabilities.
@@
  • Better Video Tokenization. This work uses image tokenizer for videos – improving tokenization to be more compact could not only enhance video quality but also enable the processing of significantly longer videos, or more efficient training on shorter videos.
  • More Modalities. Our work paves the road for learning from more modalities sources such as audio and other long sequences.

### 14 Mar 2024 v2

# World Model on Million-Length Video And Language With Blockwise RingAttention

@@ Abstract
Current language models fall short in understanding aspects of the world not easily described in words, and struggle with complex, long-form tasks. Video sequences offer valuable temporal information absent in language and static images, making them attractive for joint modeling with language. Such models could develop a understanding of both human textual knowledge and the physical world, enabling broader AI capabilities for assisting humans. However, learning from millions of tokens of video and language sequences poses challenges due to memory constraints, computational complexity, and limited datasets. To address these challenges, we curate a large dataset of diverse videos and books, utilize the Blockwise RingAttention technique to scalably train on long sequences, and gradually increase context size from 4K to 1M tokens. This paper makes the following contributions: (a) Largest context size neural network: We train one of the largest context size transformers on long video and language sequences, setting new benchmarks in difficult retrieval tasks and long video understanding. (b) Solutions for overcoming vision-language training challenges, including using masked sequence packing for mixing different sequence lengths, loss weighting to balance language and vision, and model-generated QA dataset for long sequence chat. (c) A highly-optimized implementation with RingAttention, Blockwise Transformers, masked sequence packing, and other key features for training on millions-length multimodal sequences. (d) Fully open-sourced a family of 7B parameter models capable of processing long text documents (LWM-Text, LWM-Text-Chat) and videos (LWM, LWM-Chat) of over 1M tokens. This work paves the way for training on massive datasets of long video and language to develop understanding of both human knowledge and the multimodal world, and broader capabilities.
@@ 1 Introduction
To learn from video and language sequences, we need to train a model that is capable of processing more than millions of tokens per sequence and train it on a very large dataset. However, modeling millions of tokens is extremely difficult due to high memory cost, computational complexity, and lack of suitable datasets. Luckily, we have Blockwise RingAttention (LZA, 24; LA, 23), a technique for scaling up context size arbitrarily without approximations or overheads, allowing for scalable training on long sequences. We curated a large dataset of videos and languages from public book and video datasets, consisting of videos of diverse activities and long-form books. Considering the high compute cost, we grow context size from a smaller 4K context to a larger 1M context size gradually to reduce this cost, and this approach performs well in extending context effectively. Furthermore, we identify challenges associated with training on video and language: we discovered that training on a mixture of video, image, and text is crucial for optimal performance, due to images represent higher visual quality, videos offer sequential information, and text retains language understanding. To achieve this, we implemented an efficient masked sequence packing to effectively train with different sequence lengths, rather than standard sequence packing mechanism. Moreover, determining the right balance between image, video, and text training is crucial for cross modality understanding, and we suggest a ratio that proved effective. Furthermore, to address the lack of long-form chat datasets, we developed a model-generated question-answering (QA) approach by using a short-context model to generate a QA dataset from books. We found this to be crucial for chat abilities over long sequences.
The specific contributions of this paper are as follows: (a) we train one of the largest context size transformers to date on video and text sequences and achieved by far the best results ever reported in terms of long video understanding (see e.g., Figure 1) and long context fact retrieval (see e.g., Figure 2). (b) We discover a branch of challenges associated with training on video and text sequences, and propose solutions for them: loss weighting to balance language and vision, masked sequence packing to effectively train with different sequence lengths, and model-generated QA dataset for long sequence chat. (c) A highly-optimized, open-source implementation with RingAttention with Blockwise Transformer, masked sequence packing and other key features for millions-length multimodal training. (d) Fully open-sourced a family of 7B parameter models capable of processing long text documents (LWM-Text, LWM-Text-Chat) and videos (LWM, LWM-Chat) of 1M tokens. Our work paves the way for training on massive datasets of long video and language, and is useful for future development of AI systems with an understanding of both human knowledge and the multimodal world, and broader capabilities.
@@ 3 Stage I: Learning Long-Context Language Models
This stage aims at first developing LWM-Text and LWM-Text-Chat, a set of long-context language models learned by training on progressively increasing sequence length data with RingAttention and Blockwise Transformer, and modifying positional encoding parameters to account for longer sequence lengths (see Section 3.1). The training steps for growing context size are shown in Section 3.2. In Section 3.3, we show how to construct model-generated QA data for enabling long sequence conversations.
@@ 3.1 Extending Context
Scalable Training on Long Documents. Training on long documents becomes prohibitively expensive due to memory constraints imposed by the quadratic complexity of computing the attention weights. In order to address these computational constraints, we use the Blockwise RingAttention (LZA, 24; LA, 23) implementation that leverages block-wise transformer with sequence parallelism to theoretically extend to an infinite context, bounded only by the number of devices available. We further fuse Blockwise RingAttention with FlashAttention (${}^{+}$, 22; RS, 21) using Pallas (${}^{+}$, 18) to optimize performance compared with using XLA compiler. In general, given a large enough tokens per device, the communication cost during Blockwise Transformer and RingAttention fully overlap with computation, and does not add any extra overhead.
@@ 5 Further Details
MFU.  We trained our models using TPUv4-1024, which is approximately equivalent to 450 A100s, with a batch size of 8M using FSDP Fac (23), Blockwise Transformer, and RingAttention for large contexts. Figure 9 shows the model FLOPS utilization (MFU) for each training stage. Blue color bars show language training and orange color bars show vision-language training. Our training achieves good MFUs even for very large context sizes.
@@ 6 Related Works
Our work is related to efforts that extend language models’ context windows to allow more tokens (CWCT, 23; ${}^{+}$, 23; ${}^{+}$, 23, inter alia), often by using novel extrapolation methods to expand pretrained positional encodings followed by finetuning the model on longer context data. Our model adopts a simple approach by gradually increasing $\theta$ in RoPE positional encodings along with the training context window sizes, and we found this method to be effective. There has been research into architectures that do not model pairwise interaction, such as sparse attention and sliding window (CGRS, 19; BPC, 20). Parallelization along the sequence dimension is studied in (LXLY, 21), but is not optimized for blockwise transformers and is incompatible with memory-efficient attention, which are crucial to large context training. (${}^{+}$, 22) propose hybrid usage of tensor parallelism and sequence parallelism to better reduce memory cost on a single device. Our work utilizes RingAttention and Blockwise Transformer (LZA, 24; LA, 23) to model exact pairwise interactions in very long sequences for optimal performance. Further training performance improvement is also possible with the load balancing of skipping causal masked computation (${}^{+}$, 23; ${}^{+}$, 23).
@@
Our work is also related to efforts combining vision and language (LLWL, 23; ${}^{+}$23a, ; ${}^{+}$, 23; ${}^{+}$, 23; ${}^{+}$, 23; ${}^{+}$, 23, inter alia). These efforts often use CLIP (${}^{+}$, 21) or BLIP (LLXH, 22) to encode visual information into embeddings for inputting into language models. They have the potential advantages of leveraging CLIP’s cross-modal understanding for encoding textual information in images. However, they can only be trained to predict text given visual input but not vice versa, limiting their ability to learn from diverse formats of visual and language information. In cases where they do predict visual tokens ${}^{+}$ (23), a stronger diffusion decoder is generally required due to the relatively lossy nature of CLIP embeddings. Our work, on the other hand, is autoregressive "tokens in, tokens out", allowing us to flexibly model diverse forms of image-text, text-image, text-video, video-text, and purely video, image, or text formats.
@@ 7 Conclusion
In this paper, we address the challenge of learning models to understand the world better by combining language and video. We utilize RingAttention and Blockwise Transformers to scalably train on a massive dataset of long videos and books and gradually increase sequence length, from 32K to 1M tokens, to keep compute manageable. We develop masked sequence packing and loss weighting to effectively train on a diverse dataset of videos, images, and books. Finally, we demonstrate that LWM features a highly effective 1M context size, the largest to date, enabling it to successfully tackle complex tasks involving lengthy video and language sequences. We open source our optimized implementation of Blockwise RingAttention, masked sequence packing and other key features for training on millions-length sequences, as well as a 7B parameter model capable of processing over 1M multimodal tokens. We hope this work paves the way for advancing AI models with a reliable reasoning and a grounded understanding of the world and broader capabilities.
@@
  • Better Video Tokenization. This work uses image tokenizer for videos – improving tokenization to be more compact could not only enhance video quality but also enable the processing of significantly more complex and longer videos, allowing more capable world model.
  • More Modalities. Our work paves the road for learning from additional modality sources such as audio and other long sequences, allowing a deeper and more holistic understanding about the world.